# Covariance scanning for adaptively optimal change point detection in high-dimensional linear models

### v2

@@ Appendix C Additional simulation results
See Figures C.1, C.2 and C.3 for the comparative performance of McSan, QcScan and OcScan applied with the advanced optimistic search in Algorithm 1, against with the full grid search under scenarios of (M1) and (M2) in Section 5. The results indicate only minor differences in performance, with the variants using advanced optimistic searches showing a marginal but persistent advantage overall.
@@ C.4 Additional simulations results under (M1)
Additional simulation results under the rank deficient variant of scenario (M1) in Section 5 are given in Figures C.10 and C.11. Together with Figure 3, these figures demonstrate that the performance advantage of the proposed covariance scanning methods over the competitors (MOSEG and CHARCOAL) becomes more pronounced as the rank $r$ decreases.
@@ C.5 Additional simulations results under (M2)
We present further simulation results obtained under the scenario (M2) in Section 5.

### v3

@@ Appendix C Additional simulation results
See Figures C.1, C.2 and C.3 for the comparative performance of McSan, QcScan and OcScan applied with the advanced optimistic search in Algorithm 1, against with the full grid search under scenarios of (M1) and (M2) in Section 5. The results indicate only minor differences in performance, with the variants using advanced optimistic searches showing a marginal but persistent advantage overall, possibly because the change point lies exactly on the dyadic grid explored by the advanced optimistic search.
@@ C.4 Additional simulations results under (M1)
Additional simulation results under the rank deficient variant of scenario (M1) in Section 5.1 are given in Figures C.10 and C.11. Together with Figure 3, these figures demonstrate that the performance advantage of the proposed covariance scanning methods over the competitors (MOSEG and CHARCOAL) becomes more pronounced as the rank $r$ decreases.
@@ C.5 Additional simulations results under (M2)
We present further simulation results obtained under the scenario (M2) in Section 5.1.
@@
C.6 Choice of thresholds
We investigate the choice of constants in the thresholds $\zeta_{\text{\tiny M}}=\bar{c}\widehat{\sigma}_{X}\widehat{\Psi}\sqrt{\log(p\log(n))}$ and $\zeta_{\text{\tiny Q}}=c\|\widehat{\bm{\Sigma}}\|\widehat{\Psi}^{2}\sqrt{p\log\log(n)}$, defined in 9, for McScan and QcScan, respectively. Specifically, we consider model (M2) in Section 5.1 with $\nu=0$, $n=300$ and $p=200$, under varying levels of intrinsic sparsity $\mathfrak{s}=|\bm{\Sigma}^{1/2}\bm{\delta}|_{0}=|\bm{\Sigma}^{1/2}\bm{\beta}_{0}|_{0}\in\{1,4,14,53,200\}$. The simulation results are summarised in Figures C.15 and C.16. Based on these findings, we recommend the choices $\bar{c}=1.3$ and $c=0.7$.
Figure C.15: Detection frequencies of McScan in (M2) with $\nu=0$, $n=300$ and $p=200$, obtained with various threshold choices $\zeta_{\text{\tiny M}}=\bar{c}\widehat{\sigma}_{X}\widehat{\Psi}\sqrt{\log(p\log(n))}$ where $\bar{c}\in\{1.1,1.2,1.3,1.4,1.5\}$ (left to right), when a single change point is present (red, $(q,\theta)=(1,n/4)$) and no change point exists (blue, $(q,\theta)=(0,n)$). Detection frequencies are computed from 1000 repetitions. The $x$-axis, displayed on a log scale, represents the intrinsic sparsity $\mathfrak{s}=|\bm{\Sigma}^{1/2}\bm{\delta}|_{0}$. The vertical dashed lines indicate where $\mathfrak{s}=\sqrt{p\log\log(n)}$, while the horizontal dashed lines mark the $10\%$ level.
Figure C.16: Detection frequencies of QcScan in (M2) with $\nu=0$, $n=300$ and $p=200$, obtained with various threshold choices $\zeta_{\text{\tiny Q}}=c\|\widehat{\bm{\Sigma}}\|\widehat{\Psi}^{2}\sqrt{p\log\log(n)}$ where $c\in\{0.5,0.6,0.7,0.8,0.9\}$ (left to right), when a single change point is present (red, $(q,\theta)=(1,n/4)$) and no change point exists (blue, $(q,\theta)=(0,n)$). Detection frequencies are computed from 1000 repetitions. The $x$-axis, displayed on a log scale, represents the intrinsic sparsity $\mathfrak{s}=|\bm{\Sigma}^{1/2}\bm{\delta}|_{0}$. The vertical dashed lines indicate where $\mathfrak{s}=\sqrt{p\log\log(n)}$, and the horizontal dashed lines mark the $10\%$ level.